∫sin2x/cos 2 x dx

∫sin2x/cos2x dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int \frac{sin2x}{cos^{2}x}dx
cos^{2}x = t
2cosx(-sinx)dx = dt
-sin2xdx = dt
I = \int \frac{-1}{t^{2}}dt
I = \frac{1}{t} + c
I = \frac{1}{cos^{2}x} + c
I = sec^{2}x + c